A cobb douglas production function for a new company is given by F(x,y)=50x3/5y2/5. Where x represents the units of labor and y represents the units of capital. Suppose the units of labor and the capital cost is $200 and $100 each respectively and the budget constraint is $30,000. Find the maximum production level for this manufacturer.


1
Expert's answer
2021-07-08T04:43:32-0400

The production level for this manufacturer is at maximum, if MPx/MPy = x/y.

"MPx = F'(x) = 30(y\/x)^{2\/5},"

"MPy = F'(y) = 20(x\/y)^{3\/5},"

"\frac{30(y/x)^{2/5}}{20(x/y)^{3/5}} = \frac{200} {100} ,"

1.5y/x = 2,

x = 0.75y,

If the budget constraint is $30,000, then:

200×(0.75y) + 100y = 30,000,

250y = 30,000,

y = 120 units,

x = 0.75×120 = 90 units.

"F(x, y) = 50\u00d790^{3\/5}\u00d7120^{2\/5} = 5,048.8."